OpenStudy (youssef):

I don't know how to differentiate with respect to x xy+y =1

OpenStudy (bahrom7893):

use the product rule on xy: x*dy/dx+y*1 + dy/dx = 0

OpenStudy (bahrom7893):

x(dy/dx) + y + dy/dx = 0 x(dy/dx) + (dy/dx) = -y (dy/dx)(x+1) = -y

OpenStudy (bahrom7893):

dy/dx = -y/(x+1)

OpenStudy (anonymous):

easy... uv metho.. x.(dy/dx)+y.1 + dy/dx = 0 (x-1)(dy/dx)=-y dy/dx=-y/x-1

OpenStudy (bahrom7893):

it's not x-1 it's x+1

OpenStudy (youssef):

Thanks a lot

OpenStudy (bahrom7893):

np

OpenStudy (anonymous):

yeah

OpenStudy (youssef):

It is the same answer as the doctor solve